OpenStudy (anonymous):

Determine the factors of x2 − 9x + 20.

OpenStudy (anonymous):

x^2-9x +20=0 x^2 -5x -4x +20=0 x(x-5) -4(x-5)=0 (x-4) (x-5)=0 So Factors are (x-4)(x-5)

OpenStudy (anonymous):

x^2-9x=-20 x^2-9x+20=0 x^2-5x-4x+20=0 x(x-5)-4(x-5)=0 (x-4)(x-5)=0 => Factors are (x-4)(x-5)

OpenStudy (anonymous):

how come we get the same answer

OpenStudy (anonymous):

hmmmmm that amazing

OpenStudy (anonymous):

because we derived the answer for the same question :P

OpenStudy (anonymous):

@zahabaha so x=4 and x=5